# Beyond Sparsity: Quantum Block Encoding for Dense Matrices via Hierarchically Low Rank Compression

### v2

@@ IV A Sparsification Based Quantum Approach
The decomposition algorithm from Section III provides an effective method for creating a sparse representation of a matrix. However, this process yields a set of matrices, whereas a block encoding requires a single matrix. To address this, we use an extended sparsification technique to combine these matrices into a single but larger sparse matrix.
@@
To be used in a quantum algorithm, we must analyze the properties of $A_{\text{sp}}$. As shown in Theorem III.1, the total number of non-zero elements in $A_{\text{sp}}$ scales as $O(N)$. We now consider its row and column sparsity. The non-zero entries in any given row of $A_{\text{sp}}$ come from at most three types of block matrices: a diagonal block $D_{l}$, the interpolation matrices $L_{l}$ and $R_{l-1}$, and an identity matrix $I$. By construction, the interpolation matrices at level $l$ are sparse, with at most $k_{l}$ or $k_{l}+1$ non-zeros per row or column. The diagonal blocks $D_{l}$ contain near-neighbor interactions. At the finest level ($l=1$), a block interacts with its $3^{d}-1$ neighbors, so each row in $D_{1}$ has at most $3^{d}n_{1}$ non-zeros. For coarser levels ($l>1$), a block interacts with its $(6^{d}-3^{d})$ neighbors, resulting in at most $(6^{d}-3^{d})n_{l}$ non-zeros per row in $D_{l}$. Combining these bounds, the maximum number of non-zeros in any row of $A_{\text{sp}}$ is
@@
With these properties established, we can determine the cost of creating a block encoding for $A_{\text{sp}}$. Let $c_{\text{sp}}$ be the upper bound on the magnitude of elements in the original matrix $A$, and let $f$ be the parameter defined in Theorem LABEL:apth:RRQR for the strong RRQR algorithm [guEfficientAlgorithmsComputing1996]. Based on the conditions of Theorem II.1, we arrive at the following theorem:
Theorem IV.1.
@@
While this sparsification method allows us to handle the dense matrix with algorithms designed for sparse systems, it also introduces three challenges. First, the solution vector of the extended system contains not only the desired solution $x$ but also the auxiliary variables $y_{l}$ and $z_{l}$. These additional components reduce the probability of measuring $x$, which increases the overall runtime of the algorithm. Second, the complexity of quantum linear system solvers is sensitive to the condition number of the matrix. The relationship between the condition number of the extended matrix $A_{\text{sp}}$ and the original matrix $A$ is not obvious, and the transformation could potentially make it much larger. Finally, the HBS representation is an approximation of the original dense matrix, and the impact of this initial error on the accuracy of the final quantum solution needs to be analyzed. Theorem IV.2 provides quantitative bounds for each of these issues.
Theorem IV.2.
@@
Details of Theorem IV.2 is proved in Appendix LABEL:ap:quantum. Theorem IV.2 shows that the costs associated with this sparsification method are controllable. The condition number can be bounded with regularization, and the approximation error is tied to the properties of the original system. These benefits come at a quantifiable cost in the success probability of the measurement.
V A Direct Block Encoding Algorithm for Hierarchical Matrices
@@
An alternative to the extended sparsification method is to construct a block encoding of the matrix directly from its HBS approximation. Instead of forming an extended sparse system to solve a linear equation, this approach embeds the matrix itself into a unitary operator, as defined in Definition II.2. This method requires oracles that access only $O(N)$ nonzero elements, in contrast to the $O(N^{2})$ elements required by the approach in [nguyenBlockencodingDenseFullrank2022], and it works for a broader class of kernel functions. The construction follows the hierarchical structure of the matrix shown in Equation (1), proceeding from the lowest level upwards to build the encoding for the full matrix.
@@
The construction relies on two main quantum primitives: the linear combination of unitaries (LCU) and quantum matrix multiplication (QMM). See Figures LABEL:fig:lcu_circuit and LABEL:fig:qmm_circuit for their quantum circuits, respectively. We assume that block encodings for the component matrices $D_{l}$, $L_{l}$, and $R_{l}$ are given, denoted as $U^{D}_{l}$, $U^{L}_{l}$, and $U_{l}^{R}$, respectively. Each encoding uses $a$ ancilla qubits and has an error of $\varepsilon$. The subnormalization factors for the diagonal blocks $D_{l}$ are denoted $\alpha_{l}^{(D)}$, and for the interpolation matrices $L_{l},R_{l}$ they are $\alpha_{l}^{(L)}$. Based on Lemma II.1 for block encoding sparse matrices, and recalling that the elements of $D_{l}$ are drawn from $A$, we have
@@
Definition V.1 (State preparation pair).

### v3

@@ 4 A Sparsification Based Quantum Approach
The decomposition algorithm from Section 3 provides an effective method for creating a sparse representation of a matrix. However, this process yields a set of matrices, whereas a block encoding requires a single matrix. To address this, we use an extended sparsification technique to combine these matrices into a single but larger sparse matrix.
@@
To be used in a quantum algorithm, we must analyze the properties of $A_{\text{sp}}$. As shown in Theorem 3.1, the total number of nonzero elements in $A_{\text{sp}}$ scales as $O(N)$. We now consider its row and column sparsity. The nonzero entries in any given row of $A_{\text{sp}}$ come from at most three types of block matrices: a diagonal block $D_{l}$, the interpolation matrices $L_{l}$ and $R_{l-1}$, and an identity matrix $I$. By construction, the interpolation matrices at level $l$ are sparse, with at most $k_{l}$ or $k_{l}+1$ nonzeros per row or column. The diagonal blocks $D_{l}$ contain near-neighbor interactions. At the finest level ($l=1$), a block interacts with its $3^{d}-1$ neighbors, so each row in $D_{1}$ has at most $3^{d}n_{1}$ nonzeros. For coarser levels ($l>1$), a block interacts with its $(6^{d}-3^{d})$ neighbors, resulting in at most $(6^{d}-3^{d})n_{l}$ nonzeros per row in $D_{l}$. Combining these bounds, the maximum number of nonzeros in any row of $A_{\text{sp}}$ is
@@
With these properties established, we can determine the cost of creating a block encoding for $A_{\text{sp}}$. Let $c_{\text{sp}}$ be the upper bound on the magnitude of elements in the original matrix $A$, and let $f$ be the parameter defined in Theorem 3.2 for the strong RRQR algorithm [guEfficientAlgorithmsComputing1996]. Based on the conditions of Theorem 2.1, we arrive at the following theorem:
Theorem 4.1.
@@
Proof..
By Theorem 3.1, the extended matrix $A_{\rm sp}$ stores only $O(N)$ nonzeros and has bounded row/column sparsities $s_{r},s_{c}$, with entry magnitudes bounded by $c_{\rm sp}$. Applying Lemma 2.1 to $A_{\rm sp}$ yields an $(c_{\rm{sp}}\sqrt{s_{r}s_{c}},\,\cdot,\,\varepsilon)$-block-encoding with the stated oracle/query costs.
While this sparsification method allows us to handle the dense matrix with algorithms designed for sparse systems, it also introduces three challenges. First, the solution vector of the extended system contains not only the desired solution $x$ but also the auxiliary variables $y_{l}$ and $z_{l}$. These additional components reduce the probability of measuring $x$, which increases the overall runtime of the algorithm. Second, the complexity of quantum linear system solvers is sensitive to the condition number of the matrix. The relationship between the condition number of the extended matrix $A_{\text{sp}}$ and the original matrix $A$ is not obvious, and the transformation could potentially make it much larger. Finally, the HBS representation is an approximation of the original dense matrix, and the impact of this initial error on the accuracy of the final quantum solution needs to be analyzed. Theorem 4.2 provides quantitative bounds for each of these issues.
Theorem 4.2.
@@
Theorem 4.2 shows that the costs associated with this sparsification method are controllable. The condition number can be bounded with regularization, and the approximation error is tied to the properties of the original system. These benefits come at a quantifiable cost in the success probability of the measurement.
We now justify the three claims in Theorem 4.2. For readability, the argument is organized into three corresponding parts in Section 4.1.
4.1 Analysis for the Sparsification Based Quantum Algorithm
The extended sparsification approach transforms the dense matrix into a larger but sparse linear system. While this structure is amenable to known quantum solvers, the transformation itself introduces practical issues. Here, we analyze three main aspects: the impact of auxiliary variables on the measurement success probability, the condition number of the new sparse matrix, and the error resulting from the initial hierarchical approximation.
4.1.1 Suppressing Auxiliary Variable Norms
Solving the extended linear system gives the solution vector $x^{\prime}=(x^{T},y_{1}^{T},z_{1}^{T},\ldots,y_{\lambda}^{T},z_{\lambda}^{T})^{T}$. When this vector is represented as a quantum state, the desired solution $x$ is only one component of the full state $x^{\prime}$. This means that a measurement on the final state $x^{\prime}$ yields the part corresponding to $x$ with a success probability of only $\frac{\|x\|_{2}^{2}}{\|x^{\prime}\|_{2}^{2}}$. The norms of the auxiliary variables $y_{l}$ and $z_{l}$ decrease this success probability, which in turn increases the number of repetitions needed and thus the overall runtime. We propose two methods to address this problem.
The first method, which is stated in Theorem 4.2 is to rescale the linear system. We can modify the HBS factorization by introducing a scaling parameter $t$:
$A\approx D_{1}+\left(\frac{1}{t}L_{1}\right)\left[D_{2}+\left(\frac{1}{t}L_{2}\right)\left(\cdots D_{\lambda}+\left(\frac{1}{t}L_{\lambda}\right)D_{\lambda+1}(tR_{\lambda})\cdots\right)(tR_{2})\right](tR_{1}).$ (13)
This is an exact reformulation and does not introduce additional approximation error. It modifies the extended linear system so that the new auxiliary variables become $y^{\prime}_{l}=t^{l}y_{l}$ and $z^{\prime}_{l}=t^{l}z_{l}$. By choosing a parameter $0<t<1$, the norms of the auxiliary components are suppressed exponentially with the level $l$. This increases the success probability of measuring the state for $x$. Using this modification, we have
$p_{\text{succ}}=\frac{\|x\|^{2}_{2}}{\|x^{\prime}\|_{2}^{2}}=\frac{\|x\|^{2}}{\|x\|^{2}+\sum_{l}^{\lambda}(\|y^{\prime}_{l}\|^{2}+\|z^{\prime}_{l}\|^{2})}\geq\frac{\|x\|^{2}}{\|x\|^{2}+t^{2}\sum_{l}^{\lambda}(\|y_{l}\|^{2}+\|z_{l}\|^{2})}.$ (14)
As $t$ is reduced, the failure probability decreases. This establishes item (1) in Theorem 4.2.
The second method is to apply a post-processing operator that removes the auxiliary components from the solution vector $x^{\prime}$. This is possible because the relationships between $x$, $y_{l}$, and $z_{l}$ are defined by Equation (8). We can construct an invertible matrix $A^{\prime}$ that, when applied to $x^{\prime}$, produces a vector containing only the $x$ component:
$A^{\prime}\cdot x^{\prime}=\begin{pmatrix}I&&&&&&\\ R_{1}&&-I&&&&\\ &-I&D_{2}&L_{2}&&\\ &&R_{2}&\ddots&\ddots&&\\ &&&\ddots&D_{\lambda}&L_{\lambda}&\\ &&&&R_{\lambda}&&-I\\ &&&&&-I&D_{\lambda+1}\\ \end{pmatrix}\begin{pmatrix}x\\ y_{1}\\ z_{1}\\ \vdots\\ \vdots\\ y_{\lambda}\\ z_{\lambda}\\ \end{pmatrix}=\begin{pmatrix}x\\ 0\\ 0\\ \vdots\\ \vdots\\ 0\\ 0\\ \end{pmatrix}$. (15)
[…]
4.1.2 Control of Condition Number
As noted, a potential issue with the extended sparsification is that the matrix $A_{\text{sp}}$ may be ill-conditioned. We discuss two methods to address this.
One approach is to use a preconditioner. If a sparse preconditioner $M$ can be found, we can instead solve the equivalent system:
$MAx=Mb.$
A well-chosen preconditioner ensures that the condition number of $MA$ is smaller than that of $A_{\text{sp}}$. For example, incomplete LU factorization (ILU) with thresholding constructs a sparse approximate LU decomposition of a matrix. Its parameters can be tuned to balance the reduction in the condition number against the element number of $M$. Further information can be found in [doi:10.1137/1.9781421407944, wathenPreconditioning2015].
An alternative approach is to use Tikhonov regularization to directly control the condition number, as shown in Theorem 4.2. Before introducing this method, we firstly state Gershgorin circle theorem [doi:10.1137/1.9781421407944], which provides a useful bound on the spectral radius of matrix:
Lemma 4.3.
For any $A\in\mathbb{C}^{n\times n}$, its spectrum is bounded by:
${\rm spectrum}(A)\subseteq\{z\in\mathbb{C}:|z|\leq\max\limits_{i}\sum\limits_{j=1}^{n}|a_{ij}|\}.$
We can apply this result directly to $A_{\text{sp}}$. From Theorem 3.1, the entries of $A_{\text{sp}}$ can be bounded by some constant $c_{\rm sp}>0$. Since each row of $A_{\text{sp}}$ has a constant sparsity $s_{r}$, Lemma 4.3 implies that any eigenvalue $z\in{\rm spectrum}(A_{\text{sp}})$ satisfies $|z|\leq c_{\rm sp}s_{r}$. Therefore, the largest eigenvalue of $A_{\text{sp}}$ is bounded.
However, Lemma 4.3 does not provide a lower bound on the smallest eigenvalue. To manage this problem, we employ Tikhonov regularization [englRegularizationInverseProblems1996]. Instead of solving $A_{\text{sp}}x=b$ directly, we solve a modified system. For a small parameter $\alpha>0$, we solve $(A^{\dagger}_{\text{sp}}A_{\text{sp}}+\alpha I)x^{\prime}=A^{\dagger}_{\text{sp}}b$, yielding the regularized solution $x^{\prime}=(A_{\text{sp}}^{\dagger}A_{\text{sp}}+\alpha I)^{-1}A_{\text{sp}}^{\dagger}b$. Let $R_{\alpha}=(A^{\dagger}_{\text{sp}}A_{\text{sp}}+\alpha I)^{-1}A^{\dagger}_{\text{sp}}$. The following theorem guarantees that this regularized solution converges to the true solution.
Theorem 4.4 ([englRegularizationInverseProblems1996]).
Let $A:X\to Y$ be an injective compact linear operator with dense range in $Y$. Let $f\in A(X)$ and $f^{\delta}\in Y$ satisfy
$\|f^{\delta}-f\|\leq\delta<\|f^{\delta}\|$
with $\delta>0$. Then there exists a unique parameter $\alpha=\alpha(\delta)$ such that
$\|AR_{\alpha(\delta)}f^{\delta}-f^{\delta}\|=\delta$
is satisfied and
$R_{\alpha(\delta)}f^{\delta}\to A^{-1}f,\quad\delta\to 0.$
Theorem 4.4 shows that the regularized solution converges to the true solution as the error $\delta$ in the right-hand side goes to zero. Therefore with an appropriate choice of $\alpha$, we can obtain a highly accurate solution.
We now analyze the condition number of the regularized matrix $(A^{\dagger}_{\text{sp}}A_{\text{sp}}+\alpha I)$. From Lemma 4.3, we know that $\|A_{\text{sp}}\|_{2}\leq c_{\text{sp}}s_{r}$, so
$\|A_{\text{sp}}^{\dagger}A_{\text{sp}}+\alpha I\|_{2}\leq\|A^{\dagger}_{\text{sp}}\|_{2}\|A_{\text{sp}}\|_{2}+\alpha\|I\|_{2}\leq c_{\text{sp}}^{2}s_{r}^{2}+\alpha.$
For the lower bound, for any unit vector $x$, we have $x^{\dagger}(A_{\text{sp}}^{\dagger}A_{\text{sp}}+\alpha I)x\geq\alpha$, which implies that the minimum eigenvalue of $(A^{\dagger}_{\text{sp}}A_{\text{sp}}+\alpha I)$ is bounded below by $\alpha$. Consequently, the condition number is bounded by:
$\kappa(A_{\text{sp}}^{\dagger}A_{\text{sp}}+\alpha I)\leq\frac{c_{\text{sp}}^{2}s_{r}^{2}+\alpha}{\alpha}.$
Since $(A_{\text{sp}}^{\dagger}A_{\text{sp}}+\alpha I)$ is the matrix to be inverted when solving the system of equations $(A_{\text{sp}}^{\dagger}A_{\text{sp}}+\alpha I)x^{\prime}=A_{\text{sp}}^{\dagger}b$ in the Tikhonov regularization step, this bound on its condition number is sufficient for solving the system. This proves item (2) in Theorem 4.2.
4.1.3 Error from Hierarchical Matrix Approximation
Since the extended system is built from an HBS approximation of the original matrix $A$, we also analyze the error introduced by this approximation. The error in the final solution can be bounded using a standard result from numerical linear algebra [doi:10.1137/1.9781421407944].
Theorem 4.5.
Let $A$ be the original matrix and $A_{\varepsilon}$ its compressed representation. Assume that ${\|A-A_{\varepsilon}\|}/{\|A\|}\leq\varepsilon$ and the condition $\varepsilon\kappa(A)<1$ holds, which can be ensured by the compression algorithm. Let $x$ and $b$ be vectors such that $Ax=b$. Define $b_{\varepsilon}=A_{\varepsilon}x$ and $x_{\varepsilon}=A_{\varepsilon}^{-1}b$. Then the following error bounds hold:
$\frac{\|b-b_{\varepsilon}\|}{\|b\|}\leq\varepsilon\kappa(A),$
$\frac{\|x-x_{\varepsilon}\|}{\|x\|}\leq\frac{\varepsilon\kappa(A)}{1-\varepsilon\kappa(A)},$
where $\kappa(A)$ is the condition number of $A$.
Theorem 4.5 shows that if the matrix $A$ is well-conditioned, the relative error in the solution is of the order $O(\varepsilon)$. In practice, the approximation accuracy $\varepsilon$ is chosen to be sufficiently small to ensure that the condition $\varepsilon\kappa(A)<1$ holds. Item (3) in Theorem 4.2 is exactly Theorem 4.5.
5 A Direct Block Encoding Algorithm for Hierarchical Matrices
@@
An alternative to the extended sparsification method is to construct a block encoding of the matrix directly from its HBS approximation. Instead of forming an extended sparse system to solve a linear equation, this approach embeds the matrix itself into a unitary operator, as defined in Definition 2.2. This method requires oracles that access only $O(N)$ nonzero elements, in contrast to the $O(N^{2})$ elements required by the approach in [nguyenBlockencodingDenseFullrank2022], and it works for a broader class of kernel functions. The construction follows the hierarchical structure of the matrix shown in Equation (1), proceeding from the lowest level upwards to build the encoding for the full matrix.
@@
The construction relies on two main quantum primitives: the linear combination of unitaries (LCU) and quantum matrix multiplication (QMM). See Figures LABEL:fig:lcu_circuit and LABEL:fig:qmm_circuit for their quantum circuits, respectively. We assume that block encodings for the component matrices $D_{l}$, $L_{l}$, and $R_{l}$ are given, denoted as $U^{D}_{l}$, $U^{L}_{l}$, and $U_{l}^{R}$, respectively. Each encoding uses $a$ ancilla qubits and has an error of $\varepsilon$. The subnormalization factors for the diagonal blocks $D_{l}$ are denoted $\alpha_{l}^{(D)}$, and for the interpolation matrices $L_{l},R_{l}$ they are $\alpha_{l}^{(L)}$. Based on Lemma 2.1 for block encoding sparse matrices, and recalling that the elements of $D_{l}$ are drawn from $A$, we have
@@
Definition 5.1 (State preparation pair).